OpenStudy (anonymous):

Find an equation of the line satisfying the conditions given. express your answer in standard form. Parallel to 3x-y=-5 and passing through (-1,0)

OpenStudy (anonymous):

y = 3x + b 0 = 3(-1) + b b = 3 y = 3x + 3

OpenStudy (anonymous):

in standard form 3x - y = 3

OpenStudy (anonymous):

sorry its 3x - y = -3

OpenStudy (anonymous):

parallel=same slope all you have to do is put the point and slope in a point-slope equation. slope of 3x-y=5: -3x -3x -y=5-3x y=-5+3x y=3x+5 so if y=mx+b and m is the slope, then the slope of that line is 3 plug that in with the point y-0=3(x-(-1)) y=3(x+1) y=3x+3 standard is : y-3x=3 so -3x+y=3 but you can't have a negative in the front, so 3x-y=-3